Homework: Problem Set 1 Score: 0 of 1 pt P3-5 (similar to) 8 of 14 (12 complete)Y Hw Score: 78.57%, 1 1 of 14 Question Help development project that will include rental cabins project? What if Jackson Enterprises holds the property for ten years and then Future value. Jackson Enterprises has just spent $250.000 to purchase land for a future beach front propisty has not committed to the development project, but will decide in five years whether to go forward with the project lodge, and recreational facilities. Jackson or sell off the land. Real estate values increase annually at 4.5% for property in five years if it chooses not to proceed with area. For how much can Jackson Enterprises expect to sell the
Answer #1
1) Expected sale price in 5 years = 250000*1.045^5 = $    3,11,545.48
2) Expected sale price in 10 years = 250000*1.045^10 = $    3,88,242.36